OpenStudy (anonymous):

If the width of a rectangular piece of wood is x + 2 inches long and the length is three times as long with an additional inch added, what is the area of the wood?

OpenStudy (sushi121212):

If the width is x+2, and the length is 3 times the answer plus an inch then: \[x+2\times3=3x+6\] Then, just add an inch: \[3x+6+1=3x+7\] Your answer will be 3x+7 for the length.

OpenStudy (sushi121212):

For the area, you just multiply it and get: \[x+2\times3x+7=3x ^{2}+14\] I hope this helps!